obtain an estimate for each computation by rounding the numbers so that the resulting arithmetic can easily be performed by hand or in your head. Then use a calculator to perform the computation. How reasonable is your estimate when compared to the actual answer?
0.67 X 211

Answer by checkley77(12844)

0.67 X 211
estimate: .7*200=140
actual: .67*211=141.37
140/141.37=.99 or within 99% of the actual answer.